# Understanding and Benchmarking the Trustworthiness in Multimodal LLMs for Video Understanding

### v2

@@ 3.2 Evaluated VideoLLMs
To systematically assess the trustworthiness of videoLLMs, we curated a diverse set of 23 models, spanning various design paradigms, capabilities, and accessibility levels. This includes advanced closed-source models (e.g., GPT-4o [19], Gemini [42], Claude [43]) to establish performance benchmarks, and leading open-source models (e.g., Qwen2.5-VL, LLaVA-Video, MiniCPM, InternVL) to evaluate current limitations. Detailed descriptions are provided in App. A.1.
@@ A.1 Evaluated VideoLLMs
To systematically assess the trustworthiness of VideoLLMs, we curate a diverse set of 23 models that span various design paradigms, capabilities, and accessibility levels. Our selection includes advanced closed source models (e.g., GPT-4o [19], Gemini [42] series, Claude series) to benchmark performance ceilings and highlight gaps in current open-source systems. We incorporate multiple variants from well-established frameworks such as LLaVA-Video [44] and MiniCPM [45] to examine how scaling, instruction tuning, and vision-language alignment affect model behavior. Additionally, we include models that emphasize long-context reasoning (e.g., LongVA [46]), real-time captioning (e.g., LiveCC [47]), and multi-modal instruction following (e.g., TPO [2], ShareGPT4video [48]). This composition reflects the rapid diversification of VideoLLM applications and architectures, enabling a thorough and comparative analysis across tasks requiring temporal understanding, visual grounding, and safety-critical reasoning. For detailed information of the evaluated models, see Table IV.
@@ Appendix D Evaluation Details on Truthfulness
Truthfulness is essential for the real-world deployment of large foundational models. With the rapid advancement of MLLMs, particularly those processing video data, concerns over trustworthiness and reliability have intensified. The incorporation of temporal dynamics and multimodal interactions—encompassing visual, auditory, and textual modalities—introduces significant challenges to ensuring truthfulness [4, 11, 49, 50]. Effective video understanding demands not only static visual processing but also accurate modeling of temporal sequences, contextual dependencies, and cross-modal consistency. Consequently, evaluating and enhancing the truthfulness of videoLLMs has become a critical research focus.
Prior assessments of videoLLMs have largely addressed various length of video comprehension [1, 20], hallucination phenomena [4, 11], and temporal awareness [8, 7]. However, these evaluations offer limited insight into the broader issue of truthfulness. This section proposes a comprehensive evaluation of videoLLM’s truthfulness from two perspectives: perceptual and cognitive capabilities, and reliability to contextually misleading inputs. These perspectives respectively target the models’ inherent limitations and their resilience to complex external stimuli, thus facilitating a more rigorous evaluation of videoLLM accuracy.
@@ D.1.2 Temporal Perception QA
Dataset. The dataset in this task comprises 300 videos sampled from the TempCompass [8] dataset, specifically designed to test temporal perception. These videos include scenarios with clear temporal sequences, requiring the model to identify and reason about event order and timing to select the correct answer from multiple-choice options.
@@ D.1.3 Video Description
Dataset. The dataset comprises 235 videos, including 135 sampled from OpenVid-1M [23], covering 10 diverse real-world scenarios with varying lengths, and 200 from TempCompass [8]. OpenVid-1M offers a broad range of high-quality, realistic scenes, while TempCompass focuses on videos requiring precise temporal and contextual descriptions, enabling a comprehensive evaluation of the model’s descriptive capabilities.
@@ D.2.1 Events Understanding and Detection
Dataset. We sample 200 videos from the validation set of YouCook2 [24], each paired with two text descriptions (one correct and one incorrect), resulting in 400 data points. These cooking videos feature complex event sequences, such as step-by-step procedures, requiring the model to accurately identify the correct order and details of events to avoid misinterpretations.
@@ D.2.2 Hallucination in Videos
Dataset. We sample 210 videos from the VideoHallucer [4] dataset, comprising 30 videos from each of seven categories: temporal, semantic-detail, object_relation, interaction, fact-detect, external_factual, and external_non_factual. For each item, two question types are posed: a basic question evaluating core videoLLMs capabilities, and a hallucinated question containing deliberately fabricated content. This setup assesses the model’s tendency to generate hallucinations when faced with uncertain visual inputs.
Metrics. Following the setting in VideoHallucer [4], we calculate the accuracy and the bias score as the primary metrics. For the hallucination score, we calculate the overall accuracy by considering both the basic and hallucinated questions as a paired set, marking it as a hit only if both questions are answered correctly. We posit that enhancing a model’s ability to recognize and counter hallucinations should not compromise its performance on fundamental tasks.
@@
Findings. (1) Models with 72B parameters or more consistently exhibit reduced language bias and enhanced resistance to hallucinations, indicating that large-scale parameter counts provide critical capacity for cross-modal reasoning. (2) Gemini1.5Pro and Claude-4 achieve the highest hallucination-resistance scores, suggesting that their pretraining and fine-tuning strategies are more mature in video contexts. Nevertheless, open-source 72B models (such as LLaVA-Video and Qwen2.5-VL) are rapidly closing the gap. (3) Lightweight models tend to default to “yes” responses and suffer from high false-positive rates, limiting their reliability in real-world video-understanding applications.
@@ D.3.2 Takeaways
  • Diverse strengths across model families. Open-source videoLLMs often lead in specialized video reasoning (e.g., contextual and temporal QA) thanks to task-focused modules (TPO, temporal segmentation), while closed-source models generally excel at hallucination suppression through large-scale pretraining and robust fine-tuning.
  • Scale isn’t everything. Larger parameter counts tend to improve contextual reasoning and reduce language bias in hallucination tasks, but they do not guarantee better temporal perception or event-sequence understanding—architecture design and domain adaptation can outweigh sheer size.
@@
  • Hallucination resilience correlates with conservatism. Top performers in hallucination suppression (Claude, Qwen2.5-VL-72B) exhibit a “conservative” bias—error-averse defaults that favor “no” under uncertainty—resulting in low false-positive rates but requiring careful calibration to avoid over-cautious omissions.
@@ E.1.1 Video captioning for OOD videos
Dataset. The evaluation data comes from the CapEra [25] dataset, which consists of aerial videos covering 25 different event types (such as post-earthquake scenes, floods, fires, traffic accidents, etc.). Each video corresponds to five reference captions. We randomly select 250 videos from this dataset to ensure a diverse and representative evaluation set.
@@ E.1.2 Noise Videos QA
Dataset. The evaluation dataset is derived from MVBench [3], from which we randomly sample 200 videos. Frames are uniformly sampled from each video, and noise is added with a probability of 0.3. Two types of noise are applied: Gaussian noise (continuous) and salt-and-pepper noise (discrete), both at varying intensity levels to test different degrees of corruption.
@@ E.2 Temporal Understanding Robustness
Dataset. We randomly sample 200 videos from MVBench [3] and uniformly sample frames from each video. With a probability of 0.2, we perform random frame dropping or frame shuffling operations to disrupt the temporal structure while maintaining the core content.
@@ E.3.1 Adversarial Attack for Classification
Dataset. We randomly sample 100 videos from MVBench [3] and identify high-energy regions within videos as keyframes for perturbation. We apply the MI-CWA algorithm [39] to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. These perturbations are designed to maximize prediction errors through untargeted attacks."
@@ E.3.2 Adversarial Attack for Video Captioning
Setting. Video captioning serves as a key task for evaluating the overall video understanding of videoLLMs. We apply untargeted perturbations to a few selected keyframes within each video to assess whether the model can leverage contextual information to correct the injected errors and produce accurate captions despite partial perturbations. The MI-CWA [39] algorithm is applied to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. An example is shown in Figure 18.
Dataset. We randomly sample 100 videos from MVBench [3] and identify high-energy regions within videos as keyframes for untargeted perturbation.
@@ E.4.1 The Impact of Video on Sentiment Analysis
Dataset. Videos are generated using Kling text2video system, while text samples are drawn from SST-2 [26] (100 text samples). Two evaluation scenarios are created: (1) positive text paired with natural/negative videos, and (2) negative text paired with natural/positive videos. This creates deliberate sentiment conflicts between modalities.
@@ E.4.3 The Impact of Text on Video Understanding
Dataset. We sampled 100 videos from MVBench [3] and added adversarial noise to the problems in each video. The noise is divided into three categories: Spelling errors, Grammatical errors and Noise with uncontroversial symbols.
@@ F.1 NSFW Videos Description
Metrics. We utilize three evaluation metrics to assess model performance: the Refuse-to-Answer (RtA) rate, with the template shown in Figure 24, toxicity scores derived from Google’s PerspectiveAPI [^5] https://www.perspectiveapi.com/ and a harmfulness assessment conducted by an LLM (e.g., DeepSeek) to determine if outputs are harmful or violate policy. The RtA rate quantifies the model’s propensity to reject unsafe prompts, identified through keyword matching against predefined rejection templates. Toxicity is independently evaluated using PerspectiveAPI, where we define the Perspective score as 1 minus the toxicity value and the LLM score as 1 minus the harmfulness value. Although these metrics are calculated separately, higher RtA rates typically correlate with elevated Perspective score and LLM score values in the model’s outputs.
@@
Results. The results of this task are presented in Figure 25. The score range [0, 1] is divided into 10 intervals, with the distribution of scores visualized as a heatmap. Each model’s average score—computed as (Perspective Score + RtA + LLM Score) / 3—and its RtA rate are plotted as line graphs. Compared to open-source models, commercial videoLLMs generally exhibit higher RtA rates. However, their toxicity scores are not necessarily lower; for instance, MiniCPM-V-2.6-7B and Qwen2.5-VL-7B demonstrate lower toxicity than some commercial models. Among the commercial models, GPT-4o shows the lowest response toxicity, while the Claude series achieves the highest rejection rate (up to 79%). In contrast, the Gemini series tends to produce more toxic responses, potentially due to a design preference for more direct or expressive outputs. Conversely, GPT-4o and Claude appear to favor more conservative or neutral responses, resulting in lower Perspective scores.
@@
Findings. (1) Proprietary models, including GPT-4o, Claude, and Gemini, demonstrate a stronger capability to recognize and refuse NSFW content compared to open-source models, as evidenced by higher RtA rates. (2) Open-source videoLLMs, particularly those in early stages, exhibit limited awareness of NSFW concepts, with lower RtA rates and higher toxicity scores, suggesting a need for improved safety alignment. (3) For open source models, larger model does not mean better safety alignment The performance gap between proprietary and open-source models highlights the influence of diverse and well-aligned datasets, though complex contextual reasoning remains a challenge, as seen in the variability of Avg across models.
@@ F.2 NSFW Prompts Execution
Dataset. The dataset includes 100 toxic prompts uniformly sampled from the HarmBench [27] dataset, paired with content-blank videos. These prompts are designed to provoke unsafe or inappropriate responses for large language module in videoLLMs, testing the model’s ability to recognize and reject harmful instructions regardless of the accompanying visual input.
@@ F.3 Toxic Content Continues
Dataset. The dataset comprises 80 text prompts selected from the eight toxicity categories in RealToxicityPrompts [28], with each category contributing 10 prompts with the highest toxicity scores (determined by Perspective API). Each prompt is paired with two types of videos: (1) semantically unrelated videos randomly selected, and (2) videos collected or generated to be contextually relevant to the prompt. This dual pairing tests the influence of video context on the model’s output toxicity.
@@ F.6 Deepfake Identification
Dataset. We sample 100 original videos and 100 manipulated videos from the Deepfakes Detection Entire Original dataset  [30], which is designed for deepfake detection tasks, providing a comprehensive collection of video sequences that can be used to train and evaluate deep learning models for identifying manipulated media.

### v3

@@ 3.2 Evaluated VideoLLMs
To systematically assess the trustworthiness of videoLLMs, we curated a diverse set of 23 models, spanning various design paradigms, capabilities, and accessibility levels. This includes advanced closed-source models (e.g., GPT-4o [hurst2024gpt], Gemini [team2024gemini], Claude [Anthropic]) to establish performance benchmarks, and leading open-source models (e.g., Qwen2.5-VL, LLaVA-Video, MiniCPM, InternVL) to evaluate current limitations. Detailed descriptions are provided in App. A.1.
@@ 6 Conclusion
7 Acknowledgments
This paper is supported by the National Science Foundation of China Project (Nos. 62306098 and U23B2031), Fundamental Research Funds for the Central Universities (No. JZ2024HGTB0256) and the Open Project of Anhui Provincial Key Laboratory of Multimodal Cognitive Computation, Anhui University (No. MMC202412). The computation is completed on the HPC Platform of Hefei University of Technology.
@@ A.1 Evaluated VideoLLMs
To systematically assess the trustworthiness of VideoLLMs, we curate a diverse set of 23 models that span various design paradigms, capabilities, and accessibility levels. Our selection includes advanced closed source models (e.g., GPT-4o [hurst2024gpt], Gemini [team2024gemini] series, Claude series) to benchmark performance ceilings and highlight gaps in current open-source systems. We incorporate multiple variants from well-established frameworks such as LLaVA-Video [zhang2024videoinstructiontuningsynthetic] and MiniCPM [yao2024minicpm] to examine how scaling, instruction tuning, and vision-language alignment affect model behavior. Additionally, we include models that emphasize long-context reasoning (e.g., LongVA [zhang2024longva]), real-time captioning (e.g., LiveCC [livecc]), and multi-modal instruction following (e.g., TPO [li2025temporal], ShareGPT4video [chen2024sharegpt4video]). This composition reflects the rapid diversification of VideoLLM applications and architectures, enabling a thorough and comparative analysis across tasks requiring temporal understanding, visual grounding, and safety-critical reasoning. For detailed information of the evaluated models, see Table IV.
@@ Appendix D Evaluation Details on Truthfulness
Truthfulness is essential for the real-world deployment of large foundational models. With the rapid advancement of MLLMs, particularly those processing video data, concerns over trustworthiness and reliability have intensified. The incorporation of temporal dynamics and multimodal interactions—encompassing visual, auditory, and textual modalities—introduces significant challenges to ensuring truthfulness [wang2024videohallucer, gao2025exploring, huang2024online, madan2024foundation]. Effective video understanding demands not only static visual processing but also accurate modeling of temporal sequences, contextual dependencies, and cross-modal consistency. Consequently, evaluating and enhancing the truthfulness of videoLLMs has become a critical research focus.
Prior assessments of videoLLMs have largely addressed various length of video comprehension [fu2024video, wang2024lvbench], hallucination phenomena [wang2024videohallucer, gao2025exploring], and temporal awareness [liu2024tempcompass, ning2023video]. However, these evaluations offer limited insight into the broader issue of truthfulness. This section proposes a comprehensive evaluation of videoLLM’s truthfulness from two perspectives: perceptual and cognitive capabilities, and reliability to contextually misleading inputs. These perspectives respectively target the models’ inherent limitations and their resilience to complex external stimuli, thus facilitating a more rigorous evaluation of videoLLM accuracy.
@@ D.1.2 Temporal Perception QA
Dataset. The dataset in this task comprises 300 videos sampled from the TempCompass [liu2024tempcompass] dataset, specifically designed to test temporal perception. These videos include scenarios with clear temporal sequences, requiring the model to identify and reason about event order and timing to select the correct answer from multiple-choice options.
@@ D.1.3 Video Description
Dataset. The dataset comprises 235 videos, including 135 sampled from OpenVid-1M [nan2024openvid], covering 10 diverse real-world scenarios with varying lengths, and 200 from TempCompass [liu2024tempcompass]. OpenVid-1M offers a broad range of high-quality, realistic scenes, while TempCompass focuses on videos requiring precise temporal and contextual descriptions, enabling a comprehensive evaluation of the model’s descriptive capabilities.
@@ D.2.1 Events Understanding and Detection
Dataset. We sample 200 videos from the validation set of YouCook2 [ZhXuCoAAAI18], each paired with two text descriptions (one correct and one incorrect), resulting in 400 data points. These cooking videos feature complex event sequences, such as step-by-step procedures, requiring the model to accurately identify the correct order and details of events to avoid misinterpretations.
@@ D.2.2 Hallucination in Videos
Dataset. We sample 210 videos from the VideoHallucer [wang2024videohallucer] dataset, comprising 30 videos from each of seven categories: temporal, semantic-detail, object_relation, interaction, fact-detect, external_factual, and external_non_factual. For each item, two question types are posed: a basic question evaluating core videoLLMs capabilities, and a hallucinated question containing deliberately fabricated content. This setup assesses the model’s tendency to generate hallucinations when faced with uncertain visual inputs.
Metrics. Following the setting in VideoHallucer [wang2024videohallucer], we calculate the accuracy and the bias score as the primary metrics. For the hallucination score, we calculate the overall accuracy by considering both the basic and hallucinated questions as a paired set, marking it as a hit only if both questions are answered correctly. We posit that enhancing a model’s ability to recognize and counter hallucinations should not compromise its performance on fundamental tasks.
@@
Findings. (1) Models with 72B parameters or more consistently exhibit reduced language bias and enhanced resistance to hallucinations, indicating that large-scale parameter counts provide critical capacity for cross-modal reasoning. (2) Gemini1.5-Pro and Claude‑4 achieve the highest hallucination-resistance scores, suggesting that their pretraining and fine-tuning strategies are more mature in video contexts. Nevertheless, open-source 72B models (such as LLaVA-Video and Qwen2.5‑VL) are rapidly closing the gap. (3) Lightweight models tend to default to “yes” responses and suffer from high false-positive rates, limiting their reliability in real-world video-understanding applications.
@@ D.3.2 Takeaways
  • Diverse strengths across model families. Open‑source videoLLMs often lead in specialized video reasoning (e.g., contextual and temporal QA) thanks to task‑focused modules (TPO, temporal segmentation), while closed‑source models generally excel at hallucination suppression through large‑scale pretraining and robust fine‑tuning.
  • Scale isn’t everything. Larger parameter counts tend to improve contextual reasoning and reduce language bias in hallucination tasks, but they do not guarantee better temporal perception or event‑sequence understanding—architecture design and domain adaptation can outweigh sheer size.
@@
  • Hallucination resilience correlates with conservatism. Top performers in hallucination suppression (Claude, Qwen2.5‑VL‑72B) exhibit a “conservative” bias—error‑averse defaults that favor “no” under uncertainty—resulting in low false‑positive rates but requiring careful calibration to avoid over‑cautious omissions.
@@ E.1.1 Video captioning for OOD videos
Dataset. The evaluation data comes from the CapEra [bashmal2023capera] dataset, which consists of aerial videos covering 25 different event types (such as post-earthquake scenes, floods, fires, traffic accidents, etc.). Each video corresponds to five reference captions. We randomly select 250 videos from this dataset to ensure a diverse and representative evaluation set.
@@ E.1.2 Noise Videos QA
Dataset. The evaluation dataset is derived from MVBench [li2024mvbench], from which we randomly sample 200 videos. Frames are uniformly sampled from each video, and noise is added with a probability of 0.3. Two types of noise are applied: Gaussian noise (continuous) and salt-and-pepper noise (discrete), both at varying intensity levels to test different degrees of corruption.
@@ E.2 Temporal Understanding Robustness
Dataset. We randomly sample 200 videos from MVBench [li2024mvbench] and uniformly sample frames from each video. With a probability of 0.2, we perform random frame dropping or frame shuffling operations to disrupt the temporal structure while maintaining the core content.
@@ E.3.1 Adversarial Attack for Classification
Dataset. We randomly sample 100 videos from MVBench [li2024mvbench] and identify high-energy regions within videos as keyframes for perturbation. We apply the MI-CWA algorithm [chen2023rethinking] to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. These perturbations are designed to maximize prediction errors through untargeted attacks."
@@ E.3.2 Adversarial Attack for Video Captioning
Setting. Video captioning serves as a key task for evaluating the overall video understanding of videoLLMs. We apply untargeted perturbations to a few selected keyframes within each video to assess whether the model can leverage contextual information to correct the injected errors and produce accurate captions despite partial perturbations. The MI-CWA [chen2023rethinking] algorithm is applied to generate adversarial perturbations with maximum perturbation bound $\epsilon=\frac{16}{255}$ and step size $\alpha=\frac{1}{255}$. An example is shown in Figure 18.
Dataset. We randomly sample 100 videos from MVBench [li2024mvbench] and identify high-energy regions within videos as keyframes for untargeted perturbation.
@@ E.4.1 The Impact of Video on Sentiment Analysis
Dataset. Videos are generated using Kling text2video system, while text samples are drawn from SST-2 [socher-etal-2013-recursive] (100 text samples). Two evaluation scenarios are created: (1) positive text paired with natural/negative videos, and (2) negative text paired with natural/positive videos. This creates deliberate sentiment conflicts between modalities.
@@ E.4.3 The Impact of Text on Video Understanding
Dataset. We sampled 100 videos from MVBench [li2024mvbench] and added adversarial noise to the problems in each video. The noise is divided into three categories: Spelling errors, Grammatical errors and Noise with uncontroversial symbols.
@@ F.1 NSFW Videos Description
Metrics. We use the Refuse-to-Answer (RtA) rate, with the template shown in Figure 24, non-toxicity score based on Google’s PerspectiveAPI [^5] https://www.perspectiveapi.com/ (calculated as 1 - toxicity score), and comprehensive metrics extended from average score gradient and balance index. Specifically, the refusal rate quantifies the model’s tendency to reject unsafe prompts through keyword matching against predefined refusal templates. The non-toxicity score measures the civility level of generated text. The average score, defined as the arithmetic mean of these two metrics, provides a holistic reflection of the model’s overall safety performance. Furthermore, to characterize model behavior more granularly, we introduce the average score gradient ($|\Delta|$Avg) and balance index (H-mean). The former measures the magnitude of safety performance fluctuation between adjacent models, while the latter employs the harmonic mean of refusal rate and non-toxicity score to assess the coordination between the model’s "refusal capability" and "safe output generation." A higher balance index indicates that the model can both effectively reject harmful inputs and generate safe, low-toxicity descriptions.
@@
Results. The results of this task are presented in Figure 25. The scoring range [0, 1] is discretized into several intervals and visualized as a heatmap displaying the distribution of models across multidimensional safety metrics. The blue dashed line represents the refusal rate, while the black solid line indicates the average score. Overall, commercial closed-source videoLLMs demonstrate superior performance in both refusal rate and non-toxicity score, exhibiting stronger safety alignment capabilities, whereas most open-source videoLLMs still exhibit low refusal rates and suboptimal non-toxicity scores. Among the commercial models, GPT-4o shows the lowest response toxicity, while the Claude series achieves the highest rejection rate (up to 79%). In contrast, the Gemini series tends to produce more toxic responses, potentially due to a design preference for more direct or expressive outputs. Conversely, GPT-4o and Claude appear to favor more conservative or neutral responses, resulting in lower Perspective scores.
@@
Findings. (1) Commercial models exhibit robust performance and higher safety levels. The Claude series demonstrates the highest refusal rate (>70%), while GPT-4o achieves both a lower refusal rate and the highest non-toxicity score, reflecting an excellent balance between safety and openness. In contrast, although the Gemini1.5 series shows a relatively lower refusal rate, the toxicity of its generated content is slightly inferior, indicating a design philosophy more inclined toward direct output. (2) Open-source models show significant disparities and substantial variability. Open-source videoLLMs (particularly earlier models) demonstrate limited perception of harmful content, characterized by low refusal rates and high toxicity scores, necessitating improved safety alignment. (3) The balance index reveals the stability of safety alignment. Models with high balance indices (such as GPT-4o and the Claude series) maintain effective coordination between "refusal capability" and "safe output," whereas models with low balance indices often exhibit unidirectional bias—either over-refusing or over-generating. Overall, commercial models demonstrate significantly higher safety alignment levels compared to open-source models. However, notable safety disparities persist across models. This indicates that safety alignment depends not only on model scale but is also closely related to training data, alignment strategies, and refusal template design.
@@ F.2 NSFW Prompts Execution
Dataset. The dataset includes 100 toxic prompts uniformly sampled from the HarmBench [mazeika2024harmbench] dataset, paired with content-blank videos. These prompts are designed to provoke unsafe or inappropriate responses for large language module in videoLLMs, testing the model’s ability to recognize and reject harmful instructions regardless of the accompanying visual input.
@@ F.3 Toxic Content Continues
Dataset. The dataset comprises 80 text prompts selected from the eight toxicity categories in RealToxicityPrompts [gehman2020realtoxicityprompts], with each category contributing 10 prompts with the highest toxicity scores (determined by Perspective API). Each prompt is paired with two types of videos: (1) semantically unrelated videos randomly selected, and (2) videos collected or generated to be contextually relevant to the prompt. This dual pairing tests the influence of video context on the model’s output toxicity.
@@ F.6 Deepfake Identification
Dataset. We sample 100 original videos and 100 manipulated videos from the Deepfakes Detection Entire Original dataset  [dufour2019deepfakes], which is designed for deepfake detection tasks, providing a comprehensive collection of video sequences that can be used to train and evaluate deep learning models for identifying manipulated media.